MONT 103N – Analyzing Environmental Data Review/Discussion Questions on Friedman, Hot, Flat, and Crowded April 18, 2012

Notes:

a) Because I don’t want to overburden you next week when much of the work on the final projects will be going on, you are only responsible for reading Chapters 1,2,8,11,16,17, although the rest of the book is highly recommended too(!)

b) Some questions below are adapted from the discussion guide for Hot, Flat, and Crowded developed by the publisher of the hardcover edition, Farrar, Straus and Giroux.

Discussion Questions

1. Friedman begins by outlining three trends that he says capture diverse American attitudes about energy consumption, climate change, and biodiversity: the “dumb as we wanna be” approach, the “subprime nation” mentality of borrowing our way to prosperity, and the “nation building at home” optimism of innovators who want to do what is right. What are the key aspects of each? Do you recognize these in yourself, your friends, your communities (where you grew up, here at Holy Cross)?

2. One might say that this book, written in 2007, is somewhat out of date now. Would you agree? What changed very soon after this book came out? Has the Obama administration that came into office in 2009 tried to put policies like those Friedman proposes into place? How successful have those initiatives been? Have other really significant developments about the topics Friedman discusses occurred since then? Are we any closer to addressing some of the problems he highlights?

3. In this class we have looked at some of the evidence, and the mathematical and statistical techniques scientists have used, to justify claims that climate change is real. Why do you suppose that some Americans have been so little convinced by this evidence? Why is this still such a big issue (especially for the Republican party)?

4. Chapter 2 makes the distinction between “fuels from hell” and “fuels from heaven.” How is your life fueled by both categories? What would it take to transition completely to the “heavenly” side?

5. Chapters 8 and 11 are really the heart of Friedman’s argument. And a major component of what he is saying is the notion that market demands drive innovation. What would it take to transform Americans’ perceptions so that the “Code Green” message is seen as a key to prosperity? Are there significant obstacles to that transformation at play in American society and politics at present? If so, what are they?

6. Chapter 16 focuses on the differences between China and the US. This almost forces one to consider the following questions: Can a democratic government successfully cope with the problems we are facing now? Is the Chinese system superior? What do you think?

7. What would be the consequences of becoming a “BANANA republic” (p. 406) be? Can we afford to let that happen? Or is the US in danger of becoming a “real” banana republic? Note: A number of the countries in Central America and the Caribbean (esp. Guatemala, Honduras, and Jamaica) were called “banana republics” during the late 19th and early-to-mid 20th century because they were ruled by corrupt “elected” officials and/or military dictatorships, under the domination of large US agricultural companies such as the Dole Fruit Company (Dole pineapples) and the United Fruit Company (owner of the Chiquita banana brand).

8. One of Friedman’s conclusions is that “it is much more important to change your leaders than your lightbulbs.” How do you think this will this play out in the 2012 Presidential election? How will you decide who to vote for? Will our probable “hot, flat, and crowded” future influence your choice?